OpenStudy (anonymous):

sony purchased a box of chocolates. She ate 20% of them and gave 25% to her friends from the remaining. Again, she ate 20% of it and gave 25% to her friends from the remaining. Then she gave 8 chocolates to a guest. Finally, she is left with 35% of the chocolates that she initially had. How many chocolates did sony eat?

OpenStudy (anonymous):

x/5-x=4x/5-->remaining 4x/5-1/4x=11x/20--> remaining

OpenStudy (anonymous):

11x/20-x/5=7x/20-->remaining 7x/20-x/4=2x/20--> remaining

OpenStudy (anonymous):

2x/20-8=35/100x

OpenStudy (anonymous):

or else 11x/20-4x/5=5x/20 remaining

OpenStudy (jack1):

sony purchased a box of chocolates. Total number of chocolates initially = x She ate 20% of them and gave 25% to her friends from the remaining. x(0.8) * (0.75) = y Again, she ate 20% of it and gave 25% to her friends from the remaining. y(0.8) * (0.75) = z Then she gave 8 chocolates to a guest. z-8 = 35% of x = (0.35)x Finally, she is left with 35% of the chocolates that she initially had a percentage of x, z = 0.36, so 36% giving away 8 chocolate = a drop of 1% of original quantity ==> 8 = 1% of x so 100% of x = 800 How many chocolates did sony eat? work through the above, using 800 as x

OpenStudy (jack1):

How many chocolates did sony eat? work through the above, using 800 as x x * 0.2 = eaten by sony + y * 0.2 = eaten by sony add the above

OpenStudy (anonymous):

how 0.36 %

OpenStudy (anonymous):

256 chocolates she ate